HUBBLE SPACE TELESCOPE DAILY REPORT #1001 PERIOD COVERED: 0800 EST Daily Status Report as of 299/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 4789 (Masses of Per-Main Sequence Binary Stars - Part 2) Fine Guidance Sensor #3 was used to make astrometric observations of three pre-main sequence binary stars GN-TAU, FF-TAU and HQ-TAU which are located in Taurus. The observations will be used to determine the apparent orbits of the stars which will make it possible to determine the masses of each system, since their distances are known. The observations were monitored by the OSS, and the data quality was assessed. No problems were reported. 1.2 Completed Two Sets of FOS/RD 5229 (RPT/Cal: Spectral Flat Fields - G918B2B Super Flat Redux) The Faint Object Spectrograph (Red Side) was used to make flat field spectral calibration observations using standard star WD0501+526. The peakup located the target successfully. The observations were reported to be okay, except for one observations. It was reported to have had a recentering event occur during the observation. 1.3 Completed Two Sets of PC 5072 (WF/PC Cycle 3 Calibration: PC Flats Wide. 1) Both the Wide Field Camera and the Planetary camera were used to make Earth Flat calibration exposures using the wide and medium band filters. Five observations were received and displayed in the OSS. Two of the observations were reported to be okay, and three were saturated. 1.4 Completed PC 5073 (WF/PC Cycle 3 calibration PC Flats Wide. 2) The Planetary Camera was used to make two sets of Earth flat calibration images using different filter combinations. These images will be used for flat field corrections of science images made with the PC. No problems were reported. 1.5 Completed Two Sets of PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. The observations were executed as scheduled, and no problems were reported. 1.6 Completed PC 4712 (WF/PC Cycle 2 Calibration: Measles Monitor) The Planetary Camera was used to make Earth Calibration Flat images to monitor of a form of contamination known as measles. The observations used filter F517N, both with and without filter F122M. The OSS monitored the observation and assessed the data quality. No problems were reported. 1.7 Completed WFC/PC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) The Wide Field Camera and the Planetary Camera were used to make internal flat calibration images. These images will be used to look for measles and daisies (i.e., different forms of contamination), and to monitor scattered light. The observations were executed as scheduled, and no problems were reported. 1.8 Completed WFC 5075 (WF/PC Cycle 3: Calibration: WFC Flats Wide. 1) Observations were made to make a standard calibration to obtain Earth flats using the Wide Field Camera. The OSS monitored the observations and assessed the data quality. No anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 13 Successful : 13 2.2 FHST Updates: Scheduled : 5 Successful : 5 2.3 Significant HSTARs: None 2.4 Operations Note: A new spacecraft tape recorder management technique was initiated with the SMS-298 which started yesterday, Monday, Oct. 25, 1993. Tape recorders number one and three are being used for management of science data, and tape recorder number two is being used to manage the engineering data. All indications are that the new system is working successfully. The SMGT #14(A) test was initiated today, Tuesday Oct. 26, and will end on Thursday, Oct. 28, 1993. All operations are proceeding normally, and no anomalies have been reported. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/2 5041 (GHRS Sensitivity Monitor - Cycle 3) 3.2 Complete FOS/RD/BL 5120 (HST Imagery and Spectroscopy of the Dusty SMC H II Regions N88A) 3.3 Complete Two Sets of FOS/BL 4110 (The Physics and Massive O-Stars Different Parent Galaxies) 3.5 Complete WFC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 3.6 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.7 Complete PC 5027 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor - Part 2) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-